7+-2+-11+-20+...+-119 given the following series, write each series using summation notation, using the nth term
Answer by stanbon(75887)

7+-2+-11+-20+...+-119 given the following series, write each series using summation notation, using the nth term
----
If n = 1 , 7
If n = 2 , 7-9 = -2
If n = 3 , 7-2*9 = -11
---
Solve 7-9x = -119
-9x = -126
x = 14
-------
So, if n = 15, 7-14*9 = -119
-------
Ans: summation from n = 1 to 14 of (7-(n-1)9)
===============
Cheers,
Stan H.
-------------